OpenStudy (anonymous):

The regular price of a child's entry ticket to a water park is $6 less than that for an adult's. The park offers half off of all entry tickets during the off-peak season. The Jordans paid a total of $69 for 2 adult tickets and 1 child's ticket to the water park during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket. 69 = x + one-half(x - 6) What is the regular price of a child's ticket? $38 $42 $48 $66

OpenStudy (anonymous):

lets get you to do this one. 69 = x + 1/2(x-6). First step expand the bracket then group the x's together :)

OpenStudy (anonymous):

:S Im stuck

OpenStudy (anonymous):

can you expand 1/2(x-6)?

OpenStudy (anonymous):

1/2x - 3 ?

OpenStudy (anonymous):

good

OpenStudy (anonymous):

so x +1/2(x-6) is x + 1/2x -3, ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so 69 = x +1/2x -3. Now x + 1/2 x =....?

OpenStudy (anonymous):

Ugh Im stumped again..

OpenStudy (anonymous):

you are ok with 69 = x +1/2x -3. all we did was expand the bracket. :)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

now to solve this you have to collect the x's together so you need to add x and 1/2x together

OpenStudy (anonymous):

So would it equal 1x ?

OpenStudy (anonymous):

1x + 1/2x = 3/2x. so theen u have 69 = 3/2x -3. tht ok?

OpenStudy (anonymous):

This is hard!

OpenStudy (binary3i):

let x be the regular adult ticket equation 69= x + (x-6)/2 69=x+x/2 -3 69=3x/2 -3 72=3x/2 48=x so your answer is 48.

OpenStudy (anonymous):

THANK YOU!!!

OpenStudy (binary3i):

you'r welcome sweetie